OpenStudy (elonasushchik):

Line CD contains points A (4, -7) and B (4, 8). The slope of line CD is (4 points) zero undefined positive negative

OpenStudy (elonasushchik):

@AlexandervonHumboldt2

OpenStudy (alyssa_xo):

slope formula is \(m=\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

The line is vertical, so the slope is undefined.

OpenStudy (elonasushchik):

Thankx Guys!!

OpenStudy (texaschic101):

8 - (-7) / (4 - 4) = (8 + 7) / 0 = 15/0 when there is a zero in the denominator, the slope is undefined because it is a vertical line. If the 0 would have been in the numerator, slope would have been 0